Find the area of a triangle with m∠A = 23 degrees, b = 2m, and c =8m. Round to the nearest tenth is necessary.
1
Expert's answer
2012-02-09T08:22:05-0500
The area of a triangle can be found from the equation
S =
0.5*b*c*sin(A).
S = 0.5*2*8*sin(23deg) = 3.1 m^2.